Doug is 19 years older than his sister Barbara. In 3 years, he will be twice as old as Barbara. How old is each of them now?
Answer by stanbon(75887)

Doug is 19 years older than his sister Barbara. In 3 years, he will be twice as old as Barbara. How old is each of them now?
------------
Equations:
d = b + 19
d+3 = 2(b+3)
---------------------
Substitute for "d" and solve for "b":
b+19+3 = 2b+6
b = 16 years (Barbara's age)
------
d = b + 19
d = 35 years (Doug's age)
=================
Cheers,
Stan H.
=================